# A Look at Value-Based Decision-Time vs. Background Planning Methods Across Different Settings

### v2

@@ 3.2 Modern Instantiations with Neural Network Representations
In both settings, harmful simulated experiences can again prevent the deep Dyna-Q algorithm from reaching optimal or good performance on the source tasks because of the same reasons discussed in the regular RL setting. Additionally, in the adaptation setting, after the tasks switch from the source tasks to the target tasks, deep Dyna-Q can suffer more in the adaptation process, as its model will keep generating experiences that resembles the source tasks until it adapts to the target tasks, which in the meantime will lead harmful updates to the value estimator. Also, in the generalization setting, the deep MPC algorithm will perform better on the target tasks, as while being evaluated on the target tasks it will be able to improve upon its existing policy via performing online planning; improving upon its existing policy is not possible for the deep Dyna-Q algorithm as it performs planning in an offline fashion and thus requires additional interaction with the target tasks, which is not possible in the generalization settings. More concretely, in the adaptation and generalization settings, we hypothesize that:
@@ 4.1 Experiments with Simplest Instantiations
Figure 3: (a, b) The performance of OMCP and Dyna-Q in the (a) regular RL and (b) transfer learning settings. (c-k) The performance of deep MPC and deep Dyna-Q (abbreviated as MPC and Dyna-Q) in the (c-f) regular RL and (g-k) transfer learning settings. The black dashed lines indicate the performance of the optimal policy in the corresponding environment. The green and magenta dotted lines indicate the point after which the models of the decision-time and background planning algorithms become and remain as VE models, respectively. The shaded regions are confidence intervals over (a, b) 50 and (c-k) 25 runs.
@@ 4.2 Experiments with Modern Instantiations
We now perform experiments with the deep MPC and deep Dyna-Q algorithms of Zhao et al. (2021) on MiniGrid environments to empirically validate our hypotheses in Sec. 3.2. The implementation details of these algorithms can be found in App. B.2.
@@ B.2 Implementation Details of the Modern Instantiations
Additionally, we note that in our adaptation experiments, we reinitialized the replay buffers of both planning methods after the tasks switched from the source tasks to the target tasks.

### v3

@@ 3.2 Modern Instantiations with Neural Network Representations
In both settings, harmful simulated experiences can again prevent the deep Dyna-Q algorithm from reaching optimal or good performance on the source tasks because of the same reasons discussed in the regular RL setting. Additionally, in the adaptation setting, after the tasks switch from the source tasks to the target tasks, deep Dyna-Q can suffer more in the adaptation process, as its model will keep generating experiences that resembles the source tasks until it adapts to the target tasks, which in the meantime will lead harmful updates to the value estimator. Also, in the generalization setting, if the learned model of the deep MPC algorithm becomes capable of simulating at least a few time steps of the target tasks and if the learned policies of both algorithms perform similarly on the target tasks, the deep MPC algorithm will perform better on the target tasks, as while being evaluated on the target tasks it will be able to improve upon its existing policy via performing online planning; improving upon its existing policy is not possible for the deep Dyna-Q algorithm as it performs planning in an offline fashion and thus requires additional interaction with the target tasks, which is not possible in the generalization settings. More concretely, in the adaptation and generalization settings, we hypothesize that:
@@ 4.1 Experiments with Simplest Instantiations
Figure 3: (a, b) The performance of OMCP and Dyna-Q in the (a) regular RL and (b) transfer learning settings. (c-k) The performance of deep MPC, deep Dyna-Q (abbreviated as MPC and Dyna-Q) and DQN in the (c-f) regular RL and (g-k) transfer learning settings. The black dashed lines indicate the performance of the optimal policy in the corresponding environment. The green and magenta dotted lines indicate the point after which the models of the decision-time and background planning algorithms become and remain as VE models, respectively. The shaded regions are confidence intervals over (a, b) 50 and (c-k) 25 runs.
@@ 4.2 Experiments with Modern Instantiations
We now perform experiments with the deep MPC and deep Dyna-Q algorithms of Zhao et al. (2021) on MiniGrid environments to empirically validate our hypotheses in Sec. 3.2. In addition to these two algorithms, we also perform experiments with the model-free DQN algorithm (Mnih et al., 2015) to illustrate (i) the negative effects of the usage of simulated experiences in the deep Dyna-Q algorithm and (ii) the positive effects of performing online planning in the deep MPC algorithm. We note that the only difference between deep Dyna-Q and DQN is that in the DQN algorithm the value estimator is only updated with real experiences (see line 17-20 in Alg. 1). We also note that the only difference between deep MPC and DQN is that in the DQN algorithm there is no online planning component (see line 12 in Alg. 2). The implementation details of these algorithms can be found in App. B.2.
@@
To make sure that the simulated experiences are causing the problem, we also trained the DQN algorithm on these environments. Results in Fig. 3(c)-3(f), show that the DQN algorithm achieves the same performance with the deep MPC algorithm, illustrating the negative effects of using simulated experience in the updates of the value estimator in the deep Dyna-Q algorithm.
@@
To make sure that the simulated experiences are causing the problem in the deep Dyna-Q algorithm and that the online planning component is leading to a generalization performance improvement in the deep MPC algorithm, we also trained the DQN algorithm on the RDS environments. Results in Fig. 3(g)-3(h), show that the DQN algorithm achieves the same performance with the deep MPC algorithm, again illustrating the negative effects of using simulated experience in the updates of the value estimator in the deep Dyna-Q algorithm. We can also see in Fig. 3(i)-3(k) that performing online planning significantly boosts the performance of the deep MPC algorithm compared to the DQN algorithm.
@@ B.2 Implementation Details of the Modern Instantiations
In our modern instantiation experiments, we also considered the DQN algorithm [Mnih et al., 2015]. This algorithm is just a version of the deep Dyna-Q algorithm in which real experiences are used in place of the simulated experiences for updating the value estimator. Its details and hyperparameters are the same with the deep Dyna-Q algorithm.
Additionally, we note that in our adaptation experiments, we reinitialized the replay buffers of all algorithms after the tasks switched from the source tasks to the target tasks.